You apply for a $1100 loan, which you plan to repay in 5 years. You are told by the loan officer that the amount payable when the loan is due is $2700. What rate of interest, compounded bi-weekly, are you being charged? Round answers to 2 decimal places.
Answer by Boreal(15235)

compounded 26 times per year or 130 times in 5 years
2700=1100(1+(r/26))^130
divide by 1100
2.4545=(1+(r/26))^130
logs of both sides
0.38997=130 log (1+(r/26))
divide by 130
0.003=log(1+(r/26))
raise everything to the 10 power
1.00693=1+(r/26)
subtract 1 and multiply by 26 to get the rate=0.1800 or 18%
The rule of 70 would say that 70/18%, just below 4, is the doubling time of the loan in years, which makes sense in the context of this problem.